OpenStudy (studygurl14):

Help with limits? @amistre64

OpenStudy (here_to_help15):

Ill trade help with french :)

OpenStudy (studygurl14):

 So far, I've gotten \(\Delta x = \frac{2}{n}\) and \(x_1=1 + i(\frac{2}{n})\). I then did this:  Now I'm stuck and I don't know what to do. @amistre64 @freckles